OpenStudy (anonymous):

There are 10 ripe and 5 unripe bananas in a sack. Bananas are chosen one at a time, with replacment. What is the probability of choosing 2 unripe bananas and then 2 ripe bananas? The answer is 4/81 but I'm having trouble on how they got it can anyone please explain it to me and help me understand it?

OpenStudy (anonymous):

They did say. The probability of selecting an unripe banana each time is 1/3, and the probability of selecting a ripe banana each time is 2/3. But i still can't figure out how they got it.

OpenStudy (anonymous):

The chance of getting unripe: 5/15 = 1/3 The chance of getting ripe: 10/15 = 2/3 2 unripe and 2 ripe: (1/3)(1/3)(2/3)(2/3) = 4/81

OpenStudy (anonymous):

Okay thank you I feel a little slow now. I would had never gotten that with out you're help!

OpenStudy (anonymous):

You're welcome :D